OpenStudy (anonymous):

Solve the following system of equations: 2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4

OpenStudy (anonymous):

help plz

OpenStudy (xapproachesinfinity):

subtract the second from the third to create a new equation with x and y only that's first to eliminate z but you need one more equation to have x and y only so multiply the first by 4 and the third by 3 then do the subtraction one more time to another new equation then take those two new equations to solve for x and y

OpenStudy (xapproachesinfinity):

there are various ways of doing this! the one that is more helpful is the augmented matrix but there is a lot of steps changing things to get the right answer so i think that will confuse you

OpenStudy (xapproachesinfinity):

don't wait for an answer go and work what i told you and see where you stuck

OpenStudy (anonymous):

Solve the following system of equations: 2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4 (-6, -2, -1) (6, -2, -1) (-6, 2, 1) (6, 2, 1)

OpenStudy (xapproachesinfinity):

i said don't just wait for an answer, do some work